Dave has 60 coin, all dimes and nickels, worth $4.80. How many of each type of coin does he have?
Found 2 solutions by ikleyn, addingup:
Answer by ikleyn(52925)

.
Let N be the number of nickels.

Then the number of dimes is (60-N).

The value equation is

5N + 10*(60-N) = 480  cents.


Simplify and solve for N:

5N + 600 - 10N = 480  ====>  -5N = 480 - 600  ====>  -5N = - 120  ====>  N = 24.


Answer. 24 Nickels and 60-24 = 36 dimes.

Solved.

There is entire bunch of lessons on coin problems
    - Coin problems
    - More Coin problems
    - Solving coin problems without using equations
    - Kevin and Randy Muise have a jar containing coins
    - Typical coin problems from the archive
    - More complicated coin problems
    - Solving coin problems mentally by grouping without using equations
    - Santa Claus helps solving coin problem
    - OVERVIEW of lessons on coin word problems
in this site.

You will find there the lessons for all levels - from introductory to advanced,
and for all methods used - from one equation to two equations and even without equations.

Read them and become an expert in solution of coin problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Coin problems".


Answer by addingup(3677)

d+n = 60 subtract n from both sides:
d = 60-n I'll use this value for d next.
0.10d+0.05n = 4.80 substitute for d:
0.10(60-n)+0.05n = 4.80 distribute the 0.10 over the terms in parenthesis (Distributive Properties of Multiplication)
6-0.10n+0.05n = 4.80 subtract 6 from both sides and resolve n on left:
-0.05n = -1.2 divide both sides by 0.05, and remember -/- = + Use your calculator.